OpenStudy (anonymous):

ABSOLUTE VALUE 2 | 3x - 2 | =14 I got x=14/3 or x=10/3 If it's incorrect can you please walk me through it?

OpenStudy (justnick09):

okay so first do distributive property so the equation is 6x-4=14 now add 4 to both sides 6x=18 x=3

OpenStudy (the_fizicx99):

Divide both sides by 2 | 3x - 2 | = 7 Now its an absolute value so make another equation except with a -7 | 3x - 2 | = 7 and | 3x - 2 | = -7 Add both sides 2 | 3x - 2 | = 7 +2 +2 3x = 9 /3 /3 x = 3 Can you solve you solve the other? | 3x - 2 | = -7

OpenStudy (anonymous):

I got x= 5/3 ...

OpenStudy (justnick09):

how? show us ur steps

OpenStudy (anonymous):

| 3x -2 | = -7 I added 2 to both sides and got 3x =5 Then divided both sides by 3 and got x = 5/3 e.e

OpenStudy (the_fizicx99):

-7 + 2 = ?

OpenStudy (anonymous):

I'm assuming that's where I messed up >.< ummm should that be -5?

OpenStudy (the_fizicx99):

Yea, x = -5/3

OpenStudy (the_fizicx99):

Now plug the x's you found and see if they give you a true statement. Otherwise its not a solution

OpenStudy (anonymous):

Ohh... Actually, I'm not sure how to do that...

OpenStudy (justnick09):

i think your supposed to muiltiply 2 in the equation 3x-2

OpenStudy (the_fizicx99):

I found that x = 3 Plug 3 for x, 3(3) - 2 = 7 9 - 2 = 7 7 = 7 so x = 3 is a solution Try the other one

OpenStudy (anonymous):

3 (5/3) - 2 = -7 -5 - 2 = -7

OpenStudy (the_fizicx99):

Keep in mind to put the negative sign, 3(-5/3) - 2 =-7 -5 - 2 = -7

OpenStudy (anonymous):

aghhh >.< Thank you very much

OpenStudy (anonymous):

and thanks everyone else for assisting

OpenStudy (the_fizicx99):

np